### v1

@@
[…]
Yekun Ke keyekun0628@gmail.com. Independent Researcher.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhizhou Sha shazz20@mails.tsinghua.edu.cn. Tsinghua University.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at UC Berkeley.
@@ 1 Introduction
Our contribution can be summarized as follows:
  • To the best of our knowledge, this is the first work to provide DP for the Bloom filter for membership query problems.
  • We have proved from a theoretical perspective that DPBloomfilter can achieve $(\epsilon,\delta)$-DP under the random response mechanism while preserving the same running time complexity compared with the standard Bloom filter.
  • We have proved from a theoretical perspective that when the DP parameters $\epsilon$ and $\delta$ are very small, DPBloomfilter can still maintain good utility.
  • Our simulation experiments also reflect the same effect as our theoretical results. The two confirm each other.
Roadmap. Our paper is organized as follows. In Section 2, we review related literature. Section 3 presents the preliminary of Bloom Filter and Differential Privacy. In Section 4, we outline the main results of our algorithm. In Section 5, we elaborate the derivations for the closed-form distribution of the random variable $W$, where $N$ is the $1-\delta$ quantile of $W$. Section 6 contains the proof of privacy guarantees for DPBloomfilter. Section 7 presents a detailed analysis of utility guarantees for DPBloomfilter. Section 8 restates the analysis results of running time for DPBloomfilter. Section 10 elaborates on the underlying intuitions that informed the design of the DPBloomfilter. In Section 11, we conclude our paper.
@@ 2 Related Work
In Section 2.1, we introduce the mechanism and properties, as long as some variants of the Bloom filter. In Section 2.2, we discuss several principle mechanisms used in differential privacy. In Section 2.3, we show the importance of differential privacy in contemporary data mining and recommendation systems.
@@ 2.2 Differential Privacy
Differential privacy is a technique used to defend against privacy attacks, first proposed by Dwork et al. [15]. It has become one of the most popular frameworks for ensuring privacy in theoretical analysis and a wide range of application scenarios [39, 66, 57, 14, 52, 25, 37, 40, 38, 44, 21, 54, 33, 28, 67].
Gaussian mechanism [15] and Laplace mechanism [16] of DP are widely used techniques to achieve privacy budget. These two mechanisms control the amount of privacy provided by adjusting the variance of the added noise. However, these two mechanisms are very useful when the output is continuous, but they are slightly weak when the output is discrete. However, another classic way to make a data structure private is to add a random responses mechanism [56], also called a “flip coin”. Specifically, some discrete values in the data structure are flipped with a certain probability to achieve privacy [35, 36]. By controlling the probability of flipping, a given privacy budget is achieved.
@@
Over the past decade, numerous works have applied differential privacy to data structures or deep learning models. [32] applied differential privacy to graph data structure and designed differentially node-private algorithms by projecting input graphs onto bounded-degree graphs, enhancing privacy while maintaining accuracy in realistic network analyses. [61] introduced an adaptive method for directly collecting frequent terms under local differential privacy by constructing a tree, which can overcome challenges of accuracy and utility compared to existing n-gram approaches. [19] focused on applying differential privacy to classical data mining data structures, specifically decision trees, and analyzed the balance between privacy and the utility of existing methods. [68] demonstrated the integration of differential privacy into linear sketches, ensuring privacy while maintaining high performance in processing sensitive data streams. A related work [1] introduced the BLIP mechanism, which also applies the Random Flip mechanism to the Bloom Filter. Here, we outline the differences between our work and [1] as follows: ($i$) Our proposed DPBloomFilter can satisfy $(\epsilon,\delta)-DP$, while [1] only verified that BLIP mechanism can satisfy $\epsilon$-DP; ($ii$) [1] did not provide theoretical guarantees for the utility of the BLIP mechanism.
2.3 Privacy in Data Mining and Recommendation System
The preservation of privacy is increasingly vital within the realm of data mining. recommendation systems[31] and the realm of Artificial Intelligence [25, 20, 42, 12, 43, 62, 58, 60, 30, 8, 11, 9].
In data mining, various studies have emerged that concentrate on how to extract knowledge inherited in user behavior data without compromising user privacy. For instance, [59] introduced a density-based clustering technique incorporating differential privacy. [55] delved into the application of local differential privacy (LDP) to forestall privacy violations during the aggregation of user data, in addition to investigating data poisoning attacks on LDP. Moreover, [46] has managed to maintain both efficiency and availability in mining user behavior features within specific industries while also employing differential privacy. Besides, [52] proposes a novel differentially private GNN that employs a progressive training scheme and aggregation perturbation to enhance privacy while maintaining accuracy.
On the recommendation system front, it has become common practice for streaming media and advertising companies to utilize sensitive user data, including real-time geographic locations, for user recommendations. In response to these privacy concerns, [48] first introduced a differential privacy framework tailored for recommendation systems. Besides, [6] attempts to build a Recommendation with an Attribute Protection (RAP) model, which simultaneously recommends relevant items and counters private-attribute inference attacks. More recently, [65] developed a federated recommendation framework that integrates differential privacy to shield user privacy, reducing the impact of privacy protection on recommendation quality. [27] identified a heavy reliance on user data in existing recommendation systems, leaving them susceptible to privacy breaches.
Privacy remains a critical concern in recommendation systems and data mining. This area of study is ripe for further exploration, and addressing these privacy challenges will require a substantial journey ahead.
@@ 3 Preliminary
In Section 3.1, we describe the notations we use in this paper. Section 3.2 provides the formal definition of Bloom Filter. Section 3.3 presents the formal definition of Differential Privacy, followed by a discussion on its basic composition in Section 3.4.
3.1 Notations
@@ 3.2 Bloom Filter
A Bloom filter is used to represent a set $A=\{x_{1},x_{2},\dots,x_{|A|}\}$ of $|A|$ elements from a universe $U=[n]$. A Bloom filter consists of a binary array $g\in\{0,1\}^{m}$ of $m$ bits, which are initially all set to 00, and uses $k$ independent random hash functions $h_{1},\dots,h_{k}$ with range $\{0,\dots,m-1\}$. These hash functions map each element in the universe to a random number uniform over the range $\{0,\dots,m-1\}$ for mathematical convenience. The computation time per execution for all hash functions is $\mathcal{T}_{h}$. Bloom Filter supports the following operations:
  • Init$(A)$. It takes dataset A as input. For each element $x\in A$, the bits $h_{i}(x)$ of array $g$ are set to $1$ for $1\leq i\leq k$.
  • Query$(y\in[n])$. It takes an element $y$ as input. If all $h_{i}(y)$ are set to $1$, then it outputs a binary answer to indicate that $y\in A$. If not, then it outputs $y$ is not a member of $A$.
A Bloom Filter does not have false negative issues but may yield a false positive issue, where it suggests that when a query is made to check if an element is in the set but all the positions it maps to are already set to $1$ (due to previous insertions of elements of dataset $A$). Following previous literature [35, 3, 34, 41], we assume a hash function selects each array position with equal probability. Then, the false positive rate of the Bloom Filter defined above can be mathematically approximated by the formula below
$(1-e^{-\frac{k|A|}{m}})^{k}.$
@@ 3.3 Differential Privacy
Differential Privacy (DP) ensures that the output of an algorithm remains statistically similar, under neighboring datasets introduced above, thereby protecting individual privacy. Its formal definition is as follows.
@@
For a randomized algorithm $M:U\rightarrow Range(M)$ and $\epsilon,\delta\geq 0$, if for any two neighboring data $u$ and $u^{\prime}$, it holds for $\forall Z\subset Range(M)$
$\Pr[M(u)\in Z]\leq e^{\epsilon}\Pr[M(u^{\prime})\in Z]+\delta,$
then algorithm $M$ is said to satisfy $(\epsilon,\delta)$-differentially privacy. If $\delta=0$, $M$ is called $\epsilon$-differentially private.
@@
10:procedure Init($A\subset[n],k\in\mathbb{N}_{+},m\in\mathbb{N}_{+}$) $\triangleright$ Lemma 8.1
@@
27:procedure Query($y\in[n]$) $\triangleright$ Lemma 8.2, Theorem 4.1, Theorem 4.2
@@
3.4 Basic Composition of Differential Privacy
If multiple differential privacy algorithms are involved, a composition rule becomes necessary. This section presents the simplest form of composition, as stated in the following lemma.
Lemma 3.5 (Basic composition, [24]).
@@ 4.1 Privacy for DPBloomfilter
Theorem 4.1 (Privacy for Query, informal version of Theorem 6.2).
@@
Theorem 4.1 shows that our DPBloomfilter in Algorithm 1 is $(\epsilon,\delta)$-DP. Our main technique leverages the single-bit random response technique to enhance the privacy properties of the traditional Bloom filter by composition rule (Lemma 3.5).
@@ 4.2 Utility for DPBloomfilter
Theorem 4.2 (Accuracy (compare DPBloom with true-answer) for Query, informal version of Theorem 7.4).
If the following conditions hold
  • Let $z\in\{0,1\}$ denote the true answer for whether $x\in A$.
  • Let $\widehat{z}\in\{0,1\}$ denote the answer for whether $x\in A$ output by Bloom Filter.
  • Let $\alpha:=\Pr[z=0]\in[0,1]$, $t:=e^{\epsilon_{0}}/(e^{\epsilon_{0}}+1)$, and $\delta_{\mathrm{err}}>0$.
Then, we can show
@@ 4.3 Running Complexity of DPBloomfilter
It can be proved by combining Lemma 8.1 and 8.2. ∎
@@
5 Proof for $1-\delta$ Quantile

### v2

@@
DPBloomfilter: Securing Bloom Filters with Differential Privacy
Yekun Ke and Yingyu Liang and Zhizhou Sha and Zhenmei Shi and Zhao Song and Jiahao Zhang Independent Researcher.The University of Hong Kong.University of Texas at AustinUniversity of Wisconsin-Madison.magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at UC Berkeley.
@@ 1 Introduction
Our contribution can be summarized as follows: (1) To the best of our knowledge, this is the first work to provide DP for the Bloom filter for membership query problems. (2) We have proved from a theoretical perspective that DPBloomfilter can achieve $(\epsilon,\delta)$-DP under the random response mechanism while preserving the same running time complexity compared with the standard Bloom filter. (3) We have proved from a theoretical perspective that when the DP parameters $\epsilon$ and $\delta$ are very small, DPBloomfilter can still maintain good utility. (4) Our simulation experiments also reflect the same effect as our theoretical results. The two confirm each other.
@@ 2.2 Differential Privacy
Differential privacy is a technique used to defend against privacy attacks, first proposed by Dwork et al. [10]. It has become one of the most popular frameworks for ensuring privacy in theoretical analysis and a wide range of application scenarios [30, 47, 43, 9, 39, 19, 28, 31, 29, 33, 15, 41, 24, 21, 48]. Gaussian mechanism [10] and Laplace mechanism [11] of DP are widely used techniques to achieve privacy budget. These two mechanisms control the amount of privacy provided by adjusting the variance of the added noise. However, these two mechanisms are very useful when the output is continuous, but they are slightly weak when the output is discrete. However, another classic way to make a data structure private is to add a random responses mechanism [42], also called a “flip coin”. Specifically, some discrete values in the data structure are flipped with a certain probability to achieve privacy [26, 27]. By controlling the probability of flipping, a given privacy budget is achieved. Over the past decade, numerous works have applied differential privacy to data structures or deep learning models. [23] applied differential privacy to graph data structure and designed differentially node-private algorithms by projecting input graphs onto bounded-degree graphs, enhancing privacy while maintaining accuracy in realistic network analyses. [44] introduced an adaptive method for directly collecting frequent terms under local differential privacy by constructing a tree, which can overcome challenges of accuracy and utility compared to existing n-gram approaches. [14] focused on applying differential privacy to classical data mining data structures, specifically decision trees, and analyzed the balance between privacy and the utility of existing methods. [49] demonstrated the integration of differential privacy into linear sketches, ensuring privacy while maintaining high performance in processing sensitive data streams. A related work [1] introduced the BLIP mechanism, which also applies the Random Flip mechanism to the Bloom Filter. Here, we outline the differences between our work and [1] as follows: (1) Our proposed DPBloomFilter can satisfy $(\epsilon,\delta)-DP$, while [1] only verified that BLIP mechanism can satisfy $\epsilon$-DP; (2) [1] did not provide theoretical guarantees for the utility of the BLIP mechanism.
Roadmap. Our paper is organized as follows: Section 3 presents the preliminary of Bloom Filter and Differential Privacy. In Section 4, we outline the main results of our algorithm. In Section 6, we elaborate on the underlying intuitions that informed the design of the DPBloomfilter. In Section 7, we conclude our paper.
@@ 3 Preliminary
Notations.
@@ 3.1 Bloom Filter
[…]
A Bloom Filter does not have false negative issues but may yield a false positive issue, where it suggests that when a query is made to check if an element is in the set but all the positions it maps to are already set to $1$ (due to previous insertions of elements of dataset $A$). Following previous literature [26, 3, 25, 32], we assume a hash function selects each array position with equal probability. Then, the false positive rate of the Bloom Filter defined above can be mathematically approximated by the formula as $(1-e^{-\frac{k|A|}{m}})^{k}.$
@@ 3.2 Differential Privacy
Differential Privacy (DP) ensures that the output of an algorithm remains statistically similar under neighboring datasets introduced above, thereby protecting individual privacy. Its formal definition is as follows.
@@
For a randomized algorithm $M:U\rightarrow Range(M)$ and $\epsilon,\delta\geq 0$, if for any two neighboring data $u$ and $u^{\prime}$, it holds for $\forall Z\subset Range(M),$ $\Pr[M(u)\in Z]\leq e^{\epsilon}\Pr[M(u^{\prime})\in Z]+\delta,$ then algorithm $M$ is said to satisfy $(\epsilon,\delta)$-differentially privacy. If $\delta=0$, $M$ is called $\epsilon$-differentially private.
@@
10:procedure Init($A\subset[n],k\in\mathbb{N}_{+},m\in\mathbb{N}_{+}$) $\triangleright$ Lemma E.1
@@
27:procedure Query($y\in[n]$) $\triangleright$ Lemma E.2, Theorem 4.1, Theorem 4.2
@@ 4.1 Privacy for DPBloomfilter
Theorem 4.1 (Privacy for Query, informal version of Theorem C.2).
@@
Theorem 4.1 shows that our DPBloomfilter in Algorithm 1 is $(\epsilon,\delta)$-DP. Our main technique leverages the single-bit random response technique to enhance the privacy properties of the traditional Bloom filter by composition rule (Lemma A.1).
@@ 4.2 Utility for DPBloomfilter
Theorem 4.2 (Accuracy (compare DPBloom with true-answer) for Query, informal version of Theorem D.4).
[…]
@@ 4.3 Running Complexity of DPBloomfilter
It can be proved by combining Lemma E.1 and E.2. ∎
@@
5 Technical Overview
In this section, we provide an overview of the techniques we used in proving our theoretical results.
5.1 Privacy Guarantees of Single Bit
To accomplish differential privacy, Algorithm 1 applies a random response mechanism to each bit of the standard Bloom Filter. In this section, we aim to examine the privacy guarantees for a single bit of our algorithm.
Recall that in Definition 3.1, for dataset $A\subset[n]$, we use $g[j]$ to denote the $j$-th element of array output by standard Bloom Filter. Here, we use $\widehat{g}[j]$ to denote the $j$-th element of array output by DPBloomfilter. Similarly, for any neighboring dataset $A^{\prime}\subset[n]$, we use $g^{\prime}[j]$ and $\widehat{g}^{\prime}[j]$ to denote the $j$-th element of array output by standard Bloom Filter and DPBloomfilter. To examine the privacy guarantees for the $i$-th bit, we must consider two distinct cases.
Case 1. Suppose $g^{\prime}[j]=g[j]$, then we can obtain (See also Lemma C.1) that for all $v\in\{0,1\}$, we have $\frac{\Pr[\widetilde{g}[j]=v]}{\Pr[\widetilde{g}^{\prime}[j]=v]}=1.$
Case 2. Suppose $g^{\prime}[j]\neq g[j]$, then we can obtain (See also Lemma C.1) that for all $v\in{0,1}$, we have $e^{-\epsilon_{0}}\leq\frac{\Pr[\widetilde{g}[j]=v]}{\Pr[\widetilde{g}^{\prime}[j]=v]}\leq e^{\epsilon_{0}}.$
By combining the above two cases, we can demonstrate the privacy guarantees of a single bit for our algorithm.
Lemma 5.1 (Differential Privacy for single Bit, informal version of Lemma C.1).
Let $\epsilon_{0}\geq 0$ and $\widetilde{g}[i]\in\{0,1\}$ be the $i$-th element of array output by DPBloomfilter. Then, we can show that, for all $j\in[m]$, $\widetilde{g}[j]$ is $\epsilon_{0}$-DP.
5.2 Privacy Guarantees of DPBloomFilter
Here, we comprehensively analyze the DP guarantees for our DPBloomFilter. Recall that in Definition 3.1, for dataset $A$, we use $g$ to denote the array output by standard Bloom Filter. Here, we use $\widetilde{g}$ to denote the array output by DPBloomfilter. Similarly, for any neighboring dataset $A^{\prime}$, we use $g^{\prime}$ and $\widehat{g}^{\prime}$ to denote the array output by standard Bloom Filter and DPBloomfilter, respectively. Here, we consider the set of indices $j$ within the range $m$ where the value of $g[j]$ and $g^{\prime}[j]$ differs, which is defined as $S:=\{j\in[m]:g[j]\neq g^{\prime}[j]\}.$ Thus, the set of indices $j$ where the value of $g[j]$ and $g^{\prime}[j]$ are the same can be defined as $\S.\overline{S}:=[m]\backslash$ S.
We can use the result of privacy guarantees of a single bit in Section 5.1, for any $j\in S$ and $v\in\{0,1\}$, we have $\frac{\Pr[\widetilde{g}[j]=v]}{\Pr[\widetilde{g}^{\prime}[j]=v]}=1,$ and for any $j\in\overline{S}$ and $v\in\{0,1\}$, we have
By applying the composition lemma (refer to Lemma A.1) , we obtain the following for any $Z\in\{0,1\}^{m}$,
$|\ln{\frac{\Pr[\widetilde{g}=Z]}{\Pr[\widetilde{g}^{\prime}=Z]}}|\leq|S|\epsilon_{0}.$ (1)
Here, we define $W:=|S|$ for convenience. To get a better bound for Equation 1, we need to calculate the probability distribution function of the random variable $W$. Before that, we need to define two random variables we will use. Firstly, we define $Y$ as the set of distinct values among the $k$ hash values generated by the standard Bloom filter considering one $x\in[n]$. Then we consider two data $x,x^{\prime}\in[n]$. We define $Z$ as the set of distinct values in $Y_{x}\cup Y_{x^{\prime}}$.
Then firstly we proceed to calculate the distribution of $|Y|$ (see details in Lemma B.4), we can show for any $y=1,2,\dots,k$
$\Pr[|Y|=y]=$ $~\begin{cases}1/m^{k-1},&y=1\\ \binom{m}{y}(\frac{y}{m})^{k}-\binom{m-i}{y-i}\sum_{i=1}^{k-1}\Pr[Y=i],&y=2,\cdots,k\end{cases}$
Given the probability of $|Y|$, we can calculate the conditional probability of $|Z|$ conditioned on $|Y_{x}|=a$ and $|Y_{x^{\prime}}|=b$, where $a,b\in[k]$ (see details in Lemma B.5)
$\Pr[|Z|=$ $~z||Y_{x}|=a,|Y_{x^{\prime}}|=b]$
Finally, we use the property of union probability. We can calculate the probability of $W$ (see details in Lemma B.6). Recall the notations in Section 3, we use $F_{X}^{-1}$ to denote the $1-\delta$ quantile of the Cumulative Distribution Function $F_{X}(x)$ of random variable $X$. Here, we define $N:=F_{W}^{-1}(1-\delta)$ Hence, by the properties of the quantile function, we have $\Pr[N\leq W]=1-\delta.$ By choosing the appropriate value of $\epsilon_{0}=\epsilon/N$, we have $|\ln{\frac{\Pr[\widetilde{g}=Z]}{\Pr[\widetilde{g}^{\prime}=Z]}}|\leq W\frac{\epsilon}{N}.$ Then we have, with probability $1-\delta$, $|\ln{\frac{\Pr[\widetilde{g}=Z]}{\Pr[\widetilde{g}^{\prime}=Z]}}|\leq\epsilon.$ wThen, we can demonstrate the privacy guarantees for DPBloomfilter (see also Theorem 4.1).
5.3 Utility Guarantees of DPBloomfilter
This section will present a comprehensive analysis of the utility guarantees for DPBloomfilter. We start by introducing the following conditions for the Utility guarantee of DPBloomFilter.
Condition 5.2.
We need the following conditions for Utility guarantees of DPBloomfilter:
  • Condition 1. Assume that a hash function selects each array position with equal probability.
  • Condition 2. Let $z\in\{0,1\}$ denote the ground truth for whether an element $y\in A$.
[…]
  • Condition 4. Let $\widetilde{z}\in\{0,1\}$ denote the answer output by DPBloomfilter for whether an element $y\in A$
  • Condition 5. Let $\alpha:=\Pr[z=0]\in[0,1]$
  • Condition 6. Let $t:=e^{\epsilon_{0}}/(e^{\epsilon_{0}}+1)$.
Firstly, we proceed to derive the utility of the standard Bloom Filter by calculating
$\Pr[\widehat{z}=z]=1-\Pr[\widehat{z}=1|z=0]\Pr[z=0].$
The above equation comes from the fact that Bloom Filter will not introduce a false negative. After the initialization process of Bloom Filter, the probability of one certain bit is not set to $1$ is (see also Lemma D.2) $(1-\frac{1}{m})^{|A|k}\geq e^{-2|A|k/m}.$ A false positive occurs when, for all $i\in[k]$, the elements $g[h_{i}(y)]$ are all set to $1$ after initialization. In this case, we have:
$\Pr[\widehat{z}=1|z=0]=$ $~(1-(1-\frac{1}{m})^{|A|k})^{k}$
$\leq$ $~(1-e^{-2|A|k/m})^{k}.$
Therefore, we have $\Pr[\widehat{z}=z]\geq 1-(1-e^{-2|A|k/m})^{k}\alpha.$ Further if $m=\Omega(|A|k)$ and $k=\Theta(log(\alpha/\delta_{err}))$, we have $\Pr[\widehat{z}=z]=1-\delta_{\mathrm{err}}\cdot\alpha.$
Lemma 5.3 (Accuracy for query of Standard Bloom filter, informal version of Lemma D.2).
If Condition 5.2 holds, we have $\Pr[\widehat{z}=z]\geq 1-(1-e^{-2|A|k/m})^{k}\cdot\alpha.$ Further if $m=\Omega(|A|k)$ and $k=\Theta(\log(1/\delta_{err}))$, we have $\Pr[\widehat{z}=z]\geq 1-\delta_{\mathrm{err}}\cdot\alpha.$
We then quantify the error introduced by applying the random response mechanism in the DPBloomfilter by calculating $\Pr[\widetilde{z}=\widehat{z}$]. Using basic probability rules, we have
$\Pr[\widetilde{z}=\widehat{z}$]= $~\Pr[\widetilde{z}=1|\widehat{z}=1]\Pr[\widehat{z}=1]+\Pr[\widetilde{z}=0|\widehat{z}=0]\Pr[\widehat{z}=0].$
We can compute the following term by using the definition of DPBloomfilter in Algorithm 1 (see details in Lemma D.3)
$\Pr[\widetilde{z}=1|\widehat{z}=1]=(\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1})^{k}~~~~\mathrm{and}~~~~\Pr[\widetilde{z}=0|\widehat{z}=0]\geq\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1}.$
Here we let $\Pr[\widehat{z}=0]=\widehat{\alpha}$, note that $\widehat{\alpha}=\alpha(1-\delta_{\mathrm{err}})$. Hence, $\Pr[\widehat{z}=1]=1-\Pr[\widehat{z}=0]=1-\alpha+\alpha\cdot\delta_{err}$. Then we will have (see details in Lemma D.3)
$\Pr[\widehat{z}=z]\geq t\cdot\alpha\cdot(1-\delta_{err}).$
Lemma 5.4 (Accuracy (compare DPBloomFilter with Bloom) for Query, informal version of Lemma D.3).
If Condition 5.2 holds, we can show Then, we can show $\Pr[\widetilde{z}=\widehat{z}]\geq t\cdot\alpha\cdot(1-\delta_{err}).$
Now, we can proceed to examine the utility guarantees of DPBloomfilter by calculating $\Pr[\widetilde{z}=z]$, i.e., comparing the output of DPBloomfilter with the ground truth for the query. By combining the result of the analysis above, we will have (see more details in Theorem D.4)
$\Pr[\widetilde{z}=z]\geq\alpha\cdot(1-t-t^{k})\cdot\delta_{\mathrm{err}}+\alpha\cdot t.$
Then, we demonstrated the utility guarantees of our algorithm while simultaneously ensuring privacy (see Theorem 4.2). Similar to other differential privacy algorithms, our algorithm encounters a trade-off between privacy and utility, where increased privacy typically results in a reduction in utility, and conversely. An in-depth examination of this trade-off is provided as follows.
Remark 5.5 (Trade-off between Privacy and Utility of DPBloomfilter).
An inherent trade-off exists between the privacy and utility guarantees of our algorithm. To ensure privacy, we must lower the value of $\epsilon_{0}$ in Theorem 4.1. On the other hand, for utility considerations (in Theorem 4.2), we define the lower bound of $\Pr[\widetilde{z}=z]$ as $u=\alpha(1-t-t^{k})\delta_{\mathrm{err}}+\alpha t$ , a reduction in $\epsilon_{0}$ will lead to a reduction in $t$ then finally result in a reduction in $u$. This, in turn, leads to diminished utility.
5.4 Running Time of DPBloomfilter
In this section, we will analyze the running time of our DPBloomfilter. Recall in Definition 3.1, we let $\mathcal{T}_{h}$ denote the computation time per execution for all hash functions. To analyze the algorithm’s running time, firstly, we consider the running time of initialization in Algorithm 1.
It contains two steps as follows
  • Step 1. Let’s consider the initialization of the standard Bloom Filter. For a single element $x\in A$, it needs $O(k\cdot\mathcal{T}_{h})$ time to compute over $k$ hash functions. And $|A|$ elements need to be inserted. Combining these two facts, it needs $|A|\cdot k\cdot\mathcal{T}_{h}$ time to initialize the standard Bloom Filter.
  • Step 2. Let’s consider the “Flip each bit” part in DPBloomfilter. Since there are $m$ bits in the Bloom Filter, it takes $O(m)$ time to flip each bit. Hence, it takes $O(|A|\cdot k\cdot\mathcal{T}_{h}+m)$ time to run the initialization function in Algorithm 1. (see also in Lemma E.1)
Then, we consider the running time of a single query in Algorithm 1. For each query $y$, the algorithm needs $O(k\cdot\mathcal{T}_{h})$ time to compute the hash values of $y$ over $k$ hash functions. Hence, it takes $O(k\cdot\mathcal{T}_{h})$ time to run each query $y$ in. (see also in Lemma E.2)
By combining the two running times together, we can obtain the running time of our entire algorithm is $O(|A|\cdot k\cdot\mathcal{T}_{h}+m)$. This highlights the advantage of our algorithm: it matches the time complexity of a standard Bloom Filter while providing a strong privacy guarantee.
6 Discussion
Why Random Response but not Gaussian or Laplace Noise?
As mentioned in Section 2, Gaussian and Laplace noise are two classical mechanisms to achieve differential privacy. The advantage of the Laplace mechanism is that its distribution is concentrated on its mean. Under the same privacy budget, it will not introduce too much noise like the Gaussian mechanism due to the long-tail nature of its distribution. The advantage of the Gaussian mechanism is that it has good mathematical properties and makes it easy to analyze the utility of private data structures. However, the above two mechanisms are not as effective as the random response (flip coin) mechanism when dealing with discrete values. Here, we consider the case where the discrete values are integers. Under certain privacy budgets, the noise added by Gaussian and Laplace mechanisms does not reach the threshold of $0.5$, resulting in attackers being able to remove the noise through rounding operations easily, and the privacy of the data structure no longer exists. In our case, each bit of the Bloom filter can only be $1$ or 0, which is consistent with the above situation. Hence, our work only considers the random response mechanism instead of classical Gaussian and Laplace mechanisms.
Why Flip Both 0 and $1$?
In our work, we apply a random response mechanism to each bit in the Bloom filter, either it is 0 or $1$. Although this will lead to a certain probability of false negatives in the Bloom filter, we argue that it is necessary to make the Bloom filter differentially private. Let’s consider what will happen if we don’t apply a random response mechanism like this. Suppose we only apply random responses to bits that are $1$ in the Bloom filter and leave the bits with 0 untouched. Following the notations used in Lemma A, we use $g\in\{0,1\}^{m}$ to represent the bit array generated by inserting the original dataset into the Bloom filter and $g^{\prime}\in\{0,1\}^{m}$ to represent the bit array generated by inserting the neighboring dataset into the Bloom filter. We use $\widetilde{g}$ and $\widetilde{g}^{\prime}$ to denote their private version, respectively. Without loss of generality, for some $j\in[m]$, we assume $g[j]=1$ and $g^{\prime}[j]=0$. Since we only apply a random response mechanism on bits with value $1$, then $\Pr[\widetilde{g}^{\prime}[j]=1]=0$. Therefore, we cannot calculate $\Pr[\widetilde{g}[j]=1]/\Pr[\widetilde{g}^{\prime}[j]=1]$, since the denominator is 0. Hence, we cannot have any privacy guarantees under this setting. Similar situations occur when we apply a random response mechanism on bits with value 0. We also cannot prove the differential privacy property of the Bloom filter. Therefore, we have to apply the random response mechanism on bits either with value 0 or $1$.
7 Conclusion and Future Work
In this work, we introduced DPBloomfilter, a novel approach that leverages the random response mechanism to ensure the privacy of Bloom filters. To the best of our knowledge, this is the first work that applies random response to achieve differential privacy (DP) in the membership query tasks associated with Bloom filters.
From a privacy standpoint, we have rigorously demonstrated that our method achieves $(\epsilon,\delta)$-DP while retaining the same computational complexity as the standard Bloom filter. Furthermore, our theoretical analyses, complemented by extensive experimental evaluations, confirm that the DPBloomfilter not only upholds strong privacy guarantees but also maintains high utility.
Our results open several promising avenues for future research. In particular, one interesting direction is to explore more refined trade-offs between privacy and utility, potentially by further optimizing the random response mechanism to minimize any impact on accuracy. In summary, the DPBloomfilter integrates differential privacy into the Bloom filter data structure, and we anticipate our work can advance the state-of-the-art in privacy-preserving data structures.
Appendix
Roadmap. The Appendix organizes as follows: In Section A, we introduce the notations used in the paper and differential privacy tools. In Section B, we elaborate the derivations for the closed-form distribution of the random variable $W$, where $N$ is the $1-\delta$ quantile of $W$. Section C contains the proof of privacy guarantees for DPBloomfilter. Section D presents a detailed analysis of utility guarantees for DPBloomfilter. Section E restates the analysis results of running time for DPBloomfilter.
Appendix A Basic Tools
In this section, we display the notations and basic tools for a better understanding of the readers. In Section A.1, we introduce the notations used in this paper. In Section A.2, we provide an essential basic composition Lemma for Differential Privacy.
A.1 Notations
In this section, we describe the notations we use in this paper.
A.2 Basic Composition of Differential Privacy
If multiple differential privacy algorithms are involved, a composition rule becomes necessary. This section presents the simplest form of composition, as stated as follows:
Lemma A.1 (Basic composition, [18]).
Appendix B Proof for $1-\delta$ Quantile